A step-up transformer has 23 turns on the primary coil and 800 turns on the secondary coil. If this transformer is to produce an output of 5000V with a 14-mA current, what input current and voltage are needed?

Input current = .49 A

V = ? kV
Answer #1
We know that
Vs/Vp= Ns/Np

The input voltage is
Vp = Vs[Np/Ns]
= (5000V)[(23)/(800)]
= 143.75 V
= 0.14375 kV
And Ns/Np = Ip/Is

The input current is
Ip = (Ns/Np ) Is
= ((800)/(23)) (14.0mA)
=486.9565 mA